OpenStudy (anonymous):

Please help me out. :) Find the initial velocity needed in order to shoot a projectile vertically upward a distance of 10,000 ft.

OpenStudy (amistre64):

well, we can integrate gravity to get the folrmula if need be

OpenStudy (anonymous):

am, can we use the differential calculus instead?

OpenStudy (amistre64):

acceleration is constant: a(t) = g integrate to get velocity: v(t) = gt + C, such that at t=0, C=Vi v(t) = gt + Vi integrate again to get position h(t) = 1/2 gt^2 + Vi t + C, such that t=0, C=Hi (initial height) h(t) = 1/2 gt^2 + Vi t + Hi if Hi = 0, then we reduce it to: h(t) = 1/2 gt^2 + Vi t

OpenStudy (amistre64):

not sure if i could model that with diffy Qs, havent had a lot of practice

OpenStudy (amistre64):

h(t) = 10000, and since its in feet, let g=-32

OpenStudy (anonymous):

Ouch! :( what will I do?

OpenStudy (amistre64):

10000 = -16t^2 + Vi t 0 = -16t^2 + Vi t - 10000, since the time it takes to get up to the top is the midpoint of the zeros ... its the axis of symmetry

OpenStudy (amistre64):

\[t=\frac{-V_i}{-2(16)}\] which can prolly be subbed back into the equation \[0=-16\left(\frac{-V_i}{-2(16)}\right)^2+V_i\left(\frac{-V_i}{-2(16)}\right)-10000\]

OpenStudy (amistre64):

solve for Vi ...

OpenStudy (anonymous):

Wait.

OpenStudy (anonymous):

how to solve that? :( sorry.

OpenStudy (amistre64):

your asking a differential calculus questions and cant do basic algebra?

OpenStudy (amistre64):

im pretty sure i dont need to show you how to add and multiply and cancel like terms and such ....

OpenStudy (anonymous):

sorry.. wait. sorry again.

OpenStudy (anonymous):

Vi^2 = -640000

OpenStudy (amistre64):

the negative is off, but yeah

OpenStudy (anonymous):

Vi = 800 rather. sorry.

OpenStudy (amistre64):

\[0=-16\left(\frac{-V_i}{-2(16)}\right)^2+V_i\left(\frac{-V_i}{-2(16)}\right)-10000\] \[0=-\frac{V_i^2}{2^2(16)}+\frac{V_i^2}{2(16)}-10000\] \[0=-\frac{V_i^2}{4(16)}+\frac{2V_i^2}{4(16)}-10000\] \[0=\frac{V_i^2}{4(16)}-10000\] \[10000(4)(16)=V_i^2\] \[100(2)(4)=V_i\]

OpenStudy (amistre64):

so, if it need to go 800 ft/sec just to make it to 10000ft, we need to go at least 800 or faster

OpenStudy (anonymous):

ahhh.. that's the final answer?